Reports 1-1 of 1 Clear search Modify search
AdV-SVS (Pre-commissioning)
garaventa, mehmet, sorrentino, vahlbruch - 18:04 Tuesday 11 May 2021 (51719) Print this report
New working point for CC laser

We had some indication of possible mutimode operation of the coherent control laser since a few weeks. Yesterday morning we set up the AEI source with green pump in order to observe the pump-phase error signal. The p-p error signal was about 0.2 V, see Fig 1, which is about 50% lower than in the past. By scanning the OPA length and looking at the transmitted CC beam on the SQZ_EQB1_IR_PD_MONI photodiode, we clearly identified a side resonance peak besides the carrier resonance, see Fig 2, confirming that the CC laser was running multi-mode. As we don't want to change the main laser's working point, which is already close to the Virgo main laser, we unlocked the internal PLL and changed the aux laser temperature to find a working point without multimode operation.

In terms of the DAC output to the thermal control of the laser crystal, the multi-mode operation occurs about every 13000 counts, and lasts for a temperature range of about 3000 counts. The starting point was at 32700 counts, and we found again the beat note for internal PLL at 4 MHz at the temperature 40060 counts, see Fig 3. The new working point is well far away from the multimode region, see Fig 4. When locking the OPA again, the amplitude of the pump-phase error signal is about 0.4 mV p-p, see Fig 5. The AEI source was then left in squeezing mode.

Images attached to this report
Search Help